OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (linknissan):

substitute -3 and -1 in any of the equation.. whciever gives the answer 0... that is the option@

OpenStudy (anonymous):

Hm So the formula to find equation is \[y-y1=m(x-x1)\] and  So can't you just tell?

OpenStudy (anonymous):

Plug the values.What do you think it would be? @hannah17

OpenStudy (anonymous):

Btw m=4 remember.

OpenStudy (anonymous):

y+1=4(x+3)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Please do not give answer.Guide the user to the answer. @dhiderate http://openstudy.com/code-of-conduct